# Greit-HRNet: Grouped Lightweight High-Resolution Network for Human Pose Estimation

### v1

@@ 3 Greit-HRNet
As shown in Tab. 1, our blocks have a different number of repetitions at each stage. In each branch of each stage mentioned above (except the first stage), the Greit blocks are repeated several times, as the two Greit blocks form a basic block, and the basic block is repeated several times in each branch. For the rest of this research, we use a Greit block to refer to a string of such repeated Greit blocks. For the only branch in the first stage, a Large Kernel Stem (LKS) is used for the initial processing of the feature map of the original image. As shown in Fig. 1, in other stages, a branch is assigned to a Greit block for processing feature maps, while a Greit block corresponds to one or two branches, if any. The detailed matching relationship is shown in Tab. 2. For fair comparisons with Lite-HRNet-18 [28] and Lite-HRNet-30 [28], we implement two versions of our network, Greit-HRNet-18 and Greit-HRNet-30 with different network configurations. The detailed network structure is shown in Tab. 1.
@@
As shown in Fig. 2, the LKS and Greit block follow an architecture like ShuffleNetV2 [13]. The feature map is split into two subdivisions along the channel dimension. After processing the two subdivisions to different degrees, they are concatenated with a channel shuffle operation behind them. In LKS, shown in Fig. 2(a), one subdivision first applies an LKA module [5], a depth-wise separable convolution, and a point-wise convolution, while the other subdivision expands the feature map along the channel dimension first and restores it later with a point-wise convolution. In Greit block, shown in Fig. 2(b), the processing of the feature map inside a subdivision is roughly similar to the Convolutional Block Attention Module. A GCW module is first applied to process the feature map, which is a channel attention mechanism, followed by a local sampling with a depth-wise separable convolution. A GSW module then presents a spatial attention mechanism. In addition, two adjacent branches, if any, share a GCW module. The detailed matching relationship is shown in Tab. 2.
@@ 3.1 Grouped Channel Weighting
In Lite-HRNet [28], the $X$ in Eq. 1, Eq. 2, and Eq. 3 is computed by the following two steps. First, reshape the tensors $X_{1},X_{2},\dots,X_{s-1}$ into the shape of $X_{s}$ through adaptive average pooling (AAP) operation to get the results, $X_{1}^{\prime},X_{2}^{\prime},\dots,X_{s}^{\prime}$, where $X_{s}$ denotes the tensor of the lowest resolution and $X_{i}^{\prime}$ denotes $X_{i}$ after adaptive average pooling. Second, the results, $X_{1}^{\prime},X_{2}^{\prime},\dots,X_{s}^{\prime}$ are concatenated together along the channel dimension to get $Z$. Then $Z$ is used as the $X$ in Eq. 1, Eq. 2, and Eq. 3. The advantage is that information can be exchanged among resolutions. However, in the above process, as the stage deepens, the number of channels $C_{z}$ of tensor $Z$ grows rapidly, as shown in Fig. 3. Given that the number of parameters inside $Conv_{1}$ and $Conv_{2}$ in Eq. 3 is quadratic with $C_{z}$, this processing results in a rapid increase in the overall model complexity. This is inconsistent with the former stages. In addition, a large amplitude of resolution change leads to the loss of high-resolution information at the latter stages, which is also discordant.
@@
To address these problems, we propose to use the efficient group method to replace the method mentioned above. Firstly, all branches in all stages are grouped according to Tab. 2. Specifically, these branches are divided into two groups: the high-resolution group and the low-resolution group. The high-resolution group contains branch 1 and branch 2, while the low-resolution group contains branch 3 and branch 4. Conditional channel weighting operations only take place within a single group with fewer parameters. The method can be written as,
@@
Figure 4: Global spatial information extraction. $\otimes$ denotes matrix multiplication. The first convolution sets the number of channels in the feature graph to a hyper-parameter $C^{\prime}$, and the second convolution reduces this number to 1.
@@ 3.2 Global Spatial Weighting
However, in the method mentioned above, the extraction of global context by GAP operation causes spatial information loss. Especially when the feature map is of high-resolution, the addition and average results cannot well represent the global context, as the GAP operation fails to pay attention to the pixel-level pairwise relationship and does not show preference and emphasis within the global context. Thus, we present a method to solve this problem, as shown in Fig. 4. For the tensor $X$ with scale C×H×W, where C, H, and W denote the number of channels, height, and width respectively, we first reshape it into the scale 1×C×HW to get $X^{\prime}$. Next, we perform two convolution operations on the original tensor $X$ to reduce its dimension of the channel dimension, and then reshape to get a tensor $Z$ with a scale of 1×HW×1. Finally, matrix multiplication of $X^{\prime}$ and $Z$ provides global spatial weights $Y$. This process can be written as,
@@ 4.2 Results
As shown in Tab. 3, we report the performance of our methods Greit-HRNet-18 and Greit-HRNet-30 in comparison with other state-of-the-art methods on the COCO val2017 Set. For lightweight networks, we train Greit-HRNet with input sizes 384 × 288 to achieve the highest AP score of 71.4, showing a better performance than other methods. For MobileNetV2 [17], ShuffleNetV2 [13], and EfficientHRNet-H0 [14], our model achieves better accuracy while reducing the number of model parameters by about 80$\sim$95of them. For Small HRNet [22] and LitePose-S [24], our model has fewer parameters and less computational complexity but is more accurate. In addition, our Greit-HRNet has a comparable amount of parameters and computational cost as Lite-HRNet [28] and X-HRNet [34] but has higher results than them in AP, achieving a better trade-off between model complexity and performance. For large networks, our Greit-HRNet achieves a comparable accuracy with significantly smaller model complexity compared to them.
@@
As shown in Tab. 4, we report the results on the COCO test-dev2017 Set of our Greit-HRNet and other state-of-the-art methods and demonstrate the superior efficiency of our method. Compared with Lite-HRNet, our Greit-HRNet-18 improves the AP and AR by 1.3% and 1.5% respectively, while the model and computational complexity are almost unchanged. In addition, despite a small loss of accuracy, compared with large networks, the computational cost and our model complexity are reduced by about 95%.
@@
As shown in Tab. 5, we report the results on the MPII val Set to show the superior performance of our Greit-HRNet on this dataset compared to other lightweight models. Compared to Lite-HRNet-18 and Lite-HRNet-30, our models improve by 0.7 and 0.4 points on the PCKh@0.5 score, respectively, with the same amount of model parameters and computational complexity. Compared to MobileNetV2 and ShuffleNetV2, our models have considerably fewer parameters and computational loss and achieve an accuracy of more than 2.0 4.0 points higher than them. Our Greit-HRNet-30 achieves the highest score of 87.4 PCKh@0.5, which shows the superiority of our model in optimizing the trade-off between model capacity and complexity.
@@ 4.3 Ablation Study
We conduct ablation experiments on the three modules, GCW, GSW, and LKS, and report their respective performances on the COCO val2017 set with input size 256×192, as shown in Tab. 6. To demonstrate the superior efficiency of our method, we replace conditional channel weighting modules in Lite-HRNet [28] with two stacked GCW modules to have similar model complexity and computational cost as Lite-HRNet. The overall architecture shown in Tab. 1 is used for the subsequent ablation comparison experiments.

### v2

@@ 3 Greit-HRNet
As shown in Table 1, our blocks have a different number of repetitions at each stage. In each branch of each stage mentioned above (except the first stage), the Greit blocks are repeated several times, as the two Greit blocks form a basic block, and the basic block is repeated several times in each branch. For the rest of this research, we use a Greit block to refer to a string of such repeated Greit blocks. For the only branch in the first stage, a Large Kernel Stem (LKS) is used for the initial processing of the feature map of the original image. As shown in Fig. 1, in other stages, a branch is assigned to a Greit block for processing feature maps, while a Greit block corresponds to one or two branches, if any. The detailed matching relationship is shown in Table 2. For fair comparisons with Lite-HRNet-18 [28] and Lite-HRNet-30 [28], we implement two versions of our network, Greit-HRNet-18 and Greit-HRNet-30 with different network configurations. The detailed network structure is shown in Table 1.
@@
As shown in Fig. 2, the LKS and Greit block follow an architecture like ShuffleNetV2 [13]. The feature map is split into two subdivisions along the channel dimension. After processing the two subdivisions to different degrees, they are concatenated with a channel shuffle operation behind them. In LKS, shown in Fig. 2(a), one subdivision first applies an LKA module [5], a depth-wise separable convolution, and a point-wise convolution, while the other subdivision expands the feature map along the channel dimension first and restores it later with a point-wise convolution. In Greit block, shown in Fig. 2(b), the processing of the feature map inside a subdivision is roughly similar to the Convolutional Block Attention Module. A GCW module is first applied to process the feature map, which is a channel attention mechanism, followed by a local sampling with a depth-wise separable convolution. A GSW module then presents a spatial attention mechanism. In addition, two adjacent branches, if any, share a GCW module. The detailed matching relationship is shown in Table 2.
@@ 3.1 Grouped Channel Weighting
In Lite-HRNet [28], the $X$ in Equation 1, Equation 2, and Equation 3 is computed by the following two steps. First, reshape the tensors $X_{1},X_{2},\dots,X_{s-1}$ into the shape of $X_{s}$ through adaptive average pooling (AAP) operation to get the results, $X_{1}^{\prime},X_{2}^{\prime},\dots,X_{s}^{\prime}$, where $X_{s}$ denotes the tensor of the lowest resolution and $X_{i}^{\prime}$ denotes $X_{i}$ after adaptive average pooling. Second, the results, $X_{1}^{\prime},X_{2}^{\prime},\dots,X_{s}^{\prime}$ are concatenated together along the channel dimension to get $Z$. Then $Z$ is used as the $X$ in Equation 1, Equation 2, and Equation 3. The advantage is that information can be exchanged among resolutions. However, in the above process, as the stage deepens, the number of channels $C_{z}$ of tensor $Z$ grows rapidly, as shown in Fig. 3. Given that the number of parameters inside $Conv_{1}$ and $Conv_{2}$ in Equation 3 is quadratic with $C_{z}$, this processing results in a rapid increase in the overall model complexity. This is inconsistent with the former stages. In addition, a large amplitude of resolution change leads to the loss of high-resolution information at the latter stages, which is also discordant.
@@
To address these problems, we propose to use the efficient group method to replace the method mentioned above. Firstly, all branches in all stages are grouped according to Table 2. Specifically, these branches are divided into two groups: the high-resolution group and the low-resolution group. The high-resolution group contains branch 1 and branch 2, while the low-resolution group contains branch 3 and branch 4. Conditional channel weighting operations only take place within a single group with fewer parameters. The method can be written as,
@@
(a) Information extraction
(b) Example qualitative results
Figure 4: Global spatial information extraction and example qualitative results. $\otimes$ denotes matrix multiplication. The first convolution sets the number of channels in the feature graph to a hyper-parameter $C^{\prime}$, and the second convolution reduces this number to 1.
@@ 3.2 Global Spatial Weighting
However, in the method mentioned above, the extraction of global context by GAP operation causes spatial information loss. Especially when the feature map is of high-resolution, the addition and average results cannot well represent the global context, as the GAP operation fails to pay attention to the pixel-level pairwise relationship and does not show preference and emphasis within the global context. Thus, we present a method to solve this problem, as shown in Fig. 4(a). For the tensor $X$ with scale C×H×W, where C, H, and W denote the number of channels, height, and width respectively, we first reshape it into the scale 1×C×HW to get $X^{\prime}$. Next, we perform two convolution operations on the original tensor $X$ to reduce its dimension of the channel dimension, and then reshape to get a tensor $Z$ with a scale of 1×HW×1. Finally, matrix multiplication of $X^{\prime}$ and $Z$ provides global spatial weights $Y$. This process can be written as,
@@ 4.2 Results
As shown in Table 3, we report the performance of our methods Greit-HRNet-18 and Greit-HRNet-30 in comparison with other state-of-the-art methods on the COCO val2017 Set. For lightweight networks, we train Greit-HRNet with input sizes 384 × 288 to achieve the highest AP score of 71.4, showing a better performance than other methods. For MobileNetV2 [17], ShuffleNetV2 [13], and EfficientHRNet-H0 [14], our model achieves better accuracy while reducing the number of model parameters by about 80$\sim$95of them. For Small HRNet [22] and LitePose-S [24], our model has fewer parameters and less computational complexity but is more accurate. In addition, our Greit-HRNet has a comparable amount of parameters and computational cost as Lite-HRNet [28] and X-HRNet [34] but has higher results than them in AP, achieving a better trade-off between model complexity and performance. For large networks, our Greit-HRNet achieves a comparable accuracy with significantly smaller model complexity compared to them. Fig. 4(b) shows the visual results on COCO from Greit-HRNet-30.
@@
As shown in Table 4, we report the results on the COCO test-dev2017 Set of our Greit-HRNet and other state-of-the-art methods and demonstrate the superior efficiency of our method. Compared with Lite-HRNet, our Greit-HRNet-18 improves the AP and AR by 1.3% and 1.5% respectively, while the model and computational complexity are almost unchanged. In addition, despite a small loss of accuracy, compared with large networks, the computational cost and our model complexity are reduced by about 95%.
@@
As shown in Table 5, we report the results on the MPII val Set to show the superior performance of our Greit-HRNet on this dataset compared to other lightweight models. Compared to Lite-HRNet-18 and Lite-HRNet-30, our models improve by 0.7 and 0.4 points on the PCKh@0.5 score, respectively, with the same amount of model parameters and computational complexity. Compared to MobileNetV2 and ShuffleNetV2, our models have considerably fewer parameters and computational loss and achieve an accuracy of more than 2.0 4.0 points higher than them. Our Greit-HRNet-30 achieves the highest score of 87.4 PCKh@0.5, which shows the superiority of our model in optimizing the trade-off between model capacity and complexity.
@@ 4.3 Ablation Study
We conduct ablation experiments on the three modules, GCW, GSW, and LKS, and report their respective performances on the COCO val2017 set with input size 256×192, as shown in Table 6. To demonstrate the superior efficiency of our method, we replace conditional channel weighting modules in Lite-HRNet [28] with two stacked GCW modules to have similar model complexity and computational cost as Lite-HRNet. The overall architecture shown in Table 1 is used for the subsequent ablation comparison experiments.
@@ 5 Conclusion
{credits}
5.0.1 Acknowledgements
This study was funded by Professor Yingxia Yu. We would like to express our sincere gratitude to Professor Yanxia Wang for her invaluable guidance throughout this research. We also extend our heartfelt thanks to Shuyao Shang for his assistance and support.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.